OpenStudy (anonymous):

Jewelry is commonly weighed in carats. Five carats are equivalent to one gram. How many grams of gold are contained in a 24-carat gold chain? a.120 g b.4.8 g c.5.0 g d.100 g

oregonduck (oregonduck):

5/1 = 24/x --- (5 carats to 1 gram = 24 carats to x grams) cross multiply 5x = 24 x = 24/5 x = 4 4/5 grams or b) 4.8 grams

OpenStudy (anonymous):

thankyou so much!!

oregonduck (oregonduck):

medal plz

oregonduck (oregonduck):

:)